# Gaze on the Prize: Shaping Visual Attention with Return-Guided Contrastive Learning

### v1

@@ 6 Limitations and Future Work
A key limitation of our current approach is the reliance on dense reward signals. Our return-guided contrastive learning requires sufficient return variation to generate informative training signals, but in sparse reward settings where most trajectories yield identical returns, lacks informative supervision for our contrastive learning. Future work could explore additional auxiliary signals like value estimates or curiosity rewards to provide supervision when explicit rewards are scarce burda2018exploration ; pathak2017curiosity ; badia2020never . Additionally, studies show human visual attention naturally incorporates temporal dynamics through saccades and fixations land1999roles . In contrast, our single feature based contrastive loss cannot capture motion patterns or sequential dependencies that may be critical for some complex tasks. Extending our work to address scenarios where task-relevant information spans multiple timesteps might be an exciting future direction, better mirroring the dynamic nature of biological vision systems.
@@ Appendix A Task/Environment Details
In our experiments, we evaluate seven manipulation tasks from the ManiSkill3 benchmark [33]. All tasks use single-camera RGB observations and are solvable by PPO within reasonable training time, ensuring fair comparison across methods.
@@ A.4 Cluttered Environments
To evaluate robustness to visual clutter (RQ3, Section 5.3), we create cluttered variants of four tasks by adding randomly placed distractor objects to the scene. These cluttered environments are designed to challenge the agent’s visual processing by introducing objects that are visually similar to task-relevant items but functionally irrelevant.
Appendix B Hyperparameters
The hyperparameter used in experiments are organized in Table B1.
Table B1: Hyperparameters used in experiments. All values are fixed for main experiments across tasks unless marked with *. Most hyperparameters are adopted from the original baseline implementations [33, 36] with modifications for attention and contrastive learning components.

### v2

@@ 6 Limitations and Future Work
As discussed earlier, our framework assumes that returns exhibit some degree of variation over the course the course of training. This assumption is satisfied under shaped or semi-dense reward settings commonly found in manipulation benchmarks, but becomes weak under strictly sparse-reward environments where most trajectories receive identical or near-identical returns. In such settings, the return-based ordering that drives our triplet construction provides less meaningful signal, leading the contrastive objective to collapse or become uninformative. Recent methods that employ large language models (LLMs) to automatically generate shaped rewards (e.g., Eureka ma2023eureka) offer a promising way to enrich feedback without manual engineering. However, such LLM-driven shaping does not fully resolve the fundamental challenge that sparse environments provide limited task-aligned variation for return-guided contrastive learning. A more direct approach for future work is to incorporate auxiliary signals such as curiosity-driven objectives to augment informative signals even when provided rewards are rare burda2018exploration; pathak2017curiosity; badia2020never.
Another structural constraint of our method is that our attention mechanism predicts a single fovea, modeled as a single Gaussian. While this induces a strong and structured inductive bias in certain tasks, some manipulation tasks (e.g., bimanual manipulation) require simultaneously attending to multiple spatially distinct regions within an image. An extension of our framework could predict multiple Gaussian heads (i.e., a mixture of Gaussians) with learned combination weights and adapt the same return-guided contrastive objective to supervise multiple focused regions at once. This multi-fovea variant may better capture richer spatial structure and handle tasks where multiple regions within the visual input carry task relevant information.
@@ Appendix A Task/Environment Details
In our experiments, we evaluate seven manipulation tasks from the ManiSkill3 benchmark [taomaniskill3]. All tasks use single-camera RGB observations and are solvable by PPO within reasonable training time, ensuring fair comparison across methods.
@@ A.4 Cluttered Environments
To evaluate whether return-guided contrastive learning help the agent focus on task-relevant regions in the presence of distractor objects (RQ3, Section 5.3), we create cluttered variants of four tasks by adding randomly placed distractor objects to the scene. These cluttered environments are designed to challenge the agent’s visual processing by introducing objects that are visually similar to task-relevant items but functionally irrelevant.
Appendix B Compute Overhead Analysis
To better understand the computational cost of our proposed components, we compare the baseline CNN encoder, the foveal-attention variant without contrastive learning, and the full method with both foveal attention and return-guided contrastive learning. All experiments run with 1024 parallel environments, 16-step rollouts, and a minibatch size of 512, using the same single NVIDIA RTX A6000 GPU. This controlled setup ensures a fair and reproducible comparison of compute overhead across variants.
Table B1: Compute and overhead comparison across model variants. Values are reported in milliseconds per update unless noted otherwise. Overheads for foveal attention and contrastive learning are computed relative to the total update time for each model.
Component Baseline CNN Foveal (no contrast) Full Method
Training Throughput
Steps per Second (SPS) 3820.53 2938.62 2565.78
Relative SPS 100% 76.9% (-23.1%) 67.2% (-32.8%)
Wall Time to 1M steps (seconds) 261.59 340.10 389.52
Update Cost (ms)
Total Update Time 2091.58 3247.96 3841.16
Foveal Attention Forward 1.30 1.387
Contrastive: Triplet Mining 487.44
Contrastive: FAISS Search 4.09
Contrastive: Buffer Add 0.15
Relative Overhead (%)
Foveal Attention Overhead (1.30 / 3247.96 = 0.04%) 0.04%
Contrastive Learning Overhead 12.80%
Across all variants, the majority of compute lies in the standard PPO update (forward passes through the CNN encoder, policy, and critic, along with backpropagation). Foveal attention introduces a negligible forward-pass cost (<<0.05%), while the majority of contrastive overhead arises from triplet mining, which itself contributes about 12–13% relative to the full update cost.
All per-component timings are measured using standard wall-clock timing with asynchronous GPU execution. As a result, the absolute millisecond values should be interpreted as approximate rather than precise hardware benchmarks. However, because all variants are profiled under identical conditions, the relative overhead percentages (e.g., foveal attention vs. baseline CNN vs. contrastive components) remain directly comparable and are the primary metric of interest in our compute analysis.
Appendix C CURL Implementation Details
We implement a full CURL baseline following [laskin2020curl], while replacing the original convolutional encoder with the same backbone used in our SAC agents to ensure strict apples-to-apples comparison. The agent uses $64\times 64$ RGB observations and a convolutional encoder that outputs a 256-dimensional feature shared across the actor and critic. CURL adds a bilinear contrastive head and a separate momentum encoder updated via EMA ($\tau_{\text{curl}}{=}0.005$). For each sampled batch, two independently augmented views are generated using vectorized random shifts with 4-pixel padding. The main encoder processes the query view, while the momentum encoder processes the key view, and the InfoNCE loss is computed from the resulting $B{\times}B$ similarity matrix. The CURL auxiliary loss is added to the critic update with weight $\lambda_{\text{curl}}{=}0.1$, while the actor uses encoder features detached from gradient flow. The replay buffer, training pipeline, and all SAC hyperparameters remain unchanged, ensuring that CURL differs from our main method only in the addition of the contrastive objective and momentum encoder.
Appendix D Hyperparameters
Table C1: Hyperparameters used in experiments. All values are fixed for main experiments across tasks unless marked with *. Most hyperparameters are adopted from the original baseline implementations [taomaniskill3, huang2022cleanrl] with adjustments for attention and contrastive learning components.
@@ Appendix D Hyperparameters
Top percentile 50 Sampled from high-return pool
Bottom percentile 50 Sampled from low-return pool